OpenStudy (anonymous):

What is the approximate area of the shaded region? A. 104.86 cm² B. 87.47 cm² C. 27.02 cm² D. 10.54 cm²

OpenStudy (anonymous):

here is a file

OpenStudy (anonymous):

?

OpenStudy (anonymous):

i dont get what u meant

OpenStudy (perl):

$$ \Large {Area~ of~ shaded~region= Square Area - CircleArea\\ ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~ =side^2 - \pi (radius)^2 }$$

OpenStudy (perl):

the square minus the circle will give you the remaining shaded region. imagine you bake a square cookie and you carve out a circle of the cookie. what is the remaining cookie left over

OpenStudy (anonymous):

ok thx i will medal u

OpenStudy (holly00d1248):

im srry not sure @Legends @AMYCARTER can you guys help

OpenStudy (anonymous):

Hi! So, the area of the square is equal to seven squared: 7² = 49 And the area of the circle is equal to pi times 3.5 squared and the estimate would be 38.47 rounded to the nearest hundredth: (3.14)(3.5)² ≈ 38.47 The shaded region would be the Area of Square minus the area of circle = 49 - 38.47 = 10.53 So D is an approximated answer.